OpenStudy (anonymous):

2% of x = 17 34 8.5 85 850

sam (.sam.):

\[\frac{2}{100}*x=17\]

OpenStudy (anonymous):

%=2/100 of= times

sam (.sam.):

\[x=1700/2\] x=850